# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ 2. Preliminary results
Define $\rho=\min\{1,\kappa\}$. Then for all $u\in E$ with $\|u\|=\rho$, we have
$J(u)\geq\|u\|^{\frac{n}{2}}\Big{[}\frac{2}{n}-C(\zeta\|u\|^{\tau-\frac{n}{2}}+\|u\|^{q-\frac{n}{2}})\Big{]}.$
@@
$\frac{2}{n}-C(\zeta\rho^{\tau-\frac{n}{2}}+\rho^{q-\frac{n}{2}})>0.$
@@
$J(u)\geq\rho^{\frac{n}{2}}\Big{[}\frac{2}{n}-C(\zeta\rho^{\tau-\frac{n}{2}}+\rho^{q-\frac{n}{2}})\Big{]}:=\delta>0\quad\text{for}\quad\|u\|=\rho.$
This finishes the proof of the first part of Lemma 6.3.
Next, for the proof of the second part, we assume that $u\in E\setminus\{0\}$, $u\geq 0$ with $\|u\|=1$. Define the map $\Psi\colon[1,+\infty)\to\mathbb{R}$ by $\Psi(s):=s^{-\mu}G(x,su)-G(x,u)$ for all $s\geq 1$ and $u\in\mathbb{R}^{+}:=(0,+\infty)$. By using $(g_{4})$, one can easily check that the map $\cdot\mapsto\Psi(\cdot)$ is an increasing function on $[1,+\infty)$ and thus we have $G(x,su)\geq s^{\mu}G(x,u)$ for all $s\geq 1$ and $u\in\mathbb{R}^{+}$. Now choosing $s\geq 1$ sufficiently large enough, one has
@@
we thank to $\mu>\frac{n}{2}$. The proof is now completed by choosing $e=su$ with $s$ sufficiently large. Hence the lemma is well established. ∎
@@
More precisely, we have the following crucial lemma.
@@
There exists $\lambda_{0}>0$ sufficiently large enough such that if the hypothesis $(g_{6})$ holds for all $\lambda\geq\lambda_{0}$, then the mountain pass minimax level $c$ satisfies
$0<c<c_{0}:=\min\bigg{\{}\frac{1}{2^{\frac{n}{2}}}\bigg{(}\frac{2\mu-n}{n\mu}\bigg{)}^{\frac{n}{2p}}\bigg{(}\frac{\beta_{\gamma,n}}{\alpha_{0}}\bigg{)}^{\frac{n-2}{2}},\frac{1}{2^{p}}\bigg{(}\frac{2\mu-n}{n\mu}\bigg{)}\bigg{(}\frac{\beta_{\gamma,n}}{\alpha_{0}}\bigg{)}^{\frac{(n-2)p}{n}}\bigg{\}}.$ (6.3)
@@
Let $\psi\in C^{\infty}_{0}(\mathbb{R}^{n},[0,1])$ be a cut-off function such that $\psi\equiv 1$ in $B_{\frac{1}{2}}$, $\psi\equiv 0$ in $B^{c}_{1}$, $|\nabla\psi(x)|\leq 2$ and $|\Delta\psi(x)|\leq 4$ for all $x\in\mathbb{R}^{n}$. By using $(g_{6})$, for any $s\in[0,1]$, we have
@@
Define the map $\eta\colon[0,1]\to E$ by $\eta(s)=s\psi$. Further, let us assume $\lambda_{1}>0$ be such that for all $\lambda>\lambda_{1}$, we have
@@
It follows that $\eta\in\Gamma$. Moreover, by direct calculations, one has
@@
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}<\frac{\beta_{\gamma,n}}{\alpha_{0}}.$ (6.4)
@@
By using $(g_{4})$ and the above fact, we have
$c+o_{k}(1)+o_{k}(1)\|u_{k}\|\geq\bigg{(}\frac{2}{n}-\frac{1}{\mu}\bigg{)}\Big{(}\|\Delta u_{k}\|_{p}^{p}+\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\Big{)}\quad\text{as}\quad k\to\infty.$ (6.5)
Indeed, if possible, let $\{u_{k}\}_{k}\subset E$ be unbounded in $E$. Thus we have following situations.
@@
Thus we obtain from (1.12) and (6.5) that
@@
Dividing $\|u_{k}\|^{p}$ on the both the sides and passing $k\to\infty$, we get $00\geq 2^{1-p}\big{(}\frac{2}{n}-\frac{1}{\mu}\big{)}>$ 0, which is a contradiction.
@@
$c+o_{k}(1)+o_{k}(1)\|u_{k}\|_{E}\geq 2^{1-p}\bigg{(}\frac{2}{n}-\frac{1}{\mu}\bigg{)}\|\Delta u_{k}\|_{p}^{p}\quad\text{as}\quad k\to\infty.$
Dividing $\|\Delta u_{k}\|_{p}^{p}$ on the both the sides and passing $k\to\infty$, we get $00\geq\big{(}\frac{2}{n}-\frac{1}{\mu}\big{)}>$ 0, which is again a contradiction.
@@
From the above three cases, we conclude that $\{u_{k}\}_{k}\subset E$ is a bounded sequence in $E$. Moreover, by using (6.5), one has the following estimates
$\limsup_{k\to\infty}\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\leq\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}c\quad\text{and}\quad\limsup_{k\to\infty}\|\Delta u_{k}\|_{p}^{\frac{n}{2}}\leq\bigg{(}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}c\bigg{)}^{\frac{n}{2p}}.$ (6.6)
Recalling the inequality $(a+b)^{\wp}\leq a^{\wp}+b^{\wp}$, $\forall~{}a,b\geq 0$ and $\wp\in(0,1]$, we obtain from (6.6) that
@@
Notice that $12^{\frac{2}{n-2}},\frac{n\mu}{2\mu-n}>$ 1 and $\frac{2}{n-2}<\frac{n}{(n-2)p}$. Thus one can easily deduce from the above inequality that
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}2\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}^{\frac{n}{2p}}\bigg{)}^{\frac{2}{n-2}}\Big{(}c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\Big{)}.$ (6.7)
Further, if $c\in(0,1]$, then one has $c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\leq 2c^{\frac{2}{n-2}}$. Therefore, by using (6.7), we infer that
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}2^{\frac{n}{2}}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}^{\frac{n}{2p}}c\bigg{)}^{\frac{2}{n-2}}.$ (6.8)
On the other hand, if $c\geq 1$, then we have $c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\leq 2c^{\frac{n}{(n-2)p}}$. Now again by using (6.7), we obtain
$\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}\leq\bigg{(}2^{\frac{n}{2}}\bigg{(}\frac{n\mu}{2\mu-n}\bigg{)}^{\frac{n}{2p}}c^{\frac{n}{2p}}\bigg{)}^{\frac{2}{n-2}}.$ (6.9)
In view of (6.3), (6.8) and (6.9), we conclude the proof. ∎
@@
Let the hypotheses $(g_{1})-(g_{4})$ and $(g_{6})$ be satisfied. Let $\{u_{k}\}_{k}\subset E$ be an arbitrary $(PS)_{c}$ sequence for $J$ with $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 6.4. Then there exists a subsequence of $\{u_{k}\}_{k}$ still not relabelled and $u\in E$ such that $\Delta u_{k}\to\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$. In addition, there holds
$|\Delta u_{k}|^{t-2}\Delta u_{k}\rightharpoonup|\Delta u|^{t-2}\Delta u\quad\text{in}\quad L^{\frac{t}{t-1}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty\quad\text{for}\quad t\in\{p,\frac{n}{2}\}.$
@@
Let $0<c<c_{0}$ and $\{u_{k}\}_{k}\subset E$ be a $(PS)_{c}$ sequence for $J$. By using Lemma 6.5, $\{u_{k}\}_{k}$ is a bounded sequence in $E$ and (6.4) holds. Hence we can assume without loss of generality, up to a subsequence, still not relabelled and $u\in E$ such that $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$. Now by using Lemma 2.5, Theorem 1.7 and Lemma 2.7, we have
$\begin{cases}u_{k}\to u\quad\text{in}\quad L^{\theta}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{for all}\quad\theta\in[1,+\infty);\\ u_{k}\to u\quad\text{in}\quad L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)\quad\text{for all}\quad\varrho\in[p^{\ast},+\infty);\\ \nabla u_{k}\to\nabla u\quad\text{in}\quad L^{t}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{for}\quad t\in\{p,\frac{n}{2}\};\\ u_{k}\to u\quad\text{a.e.~{}~{}in}\quad\mathbb{R}^{n}\quad\text{as}\quad k\to\infty.\end{cases}$ (6.10)
@@
$\langle{J^{\prime}(u_{k})-J^{\prime}(u),(u_{k}-u)\eta}\rangle=o_{k}(1)\quad\text{as}\quad k\to\infty.$ (6.11)
@@
for any $k\in\mathbb{N}$ and $t\in\{p,\frac{n}{2}\}$. In view of the well known Simon’s inequality with $n\geq 4$ (for instance, see to (6.27)) and (6.11), there exists $c_{\frac{n}{2}}>0$ such that
$c^{-1}_{\frac{n}{2}}\int_{B_{R}}|\Delta u_{k}-\Delta u|^{\frac{n}{2}}\,\mathrm{d}x$ $\leq\int_{B_{R}}\textit{P}^{\frac{n}{2}}_{k}(u_{k},u)\,\mathrm{d}x\leq\displaystyle\sum_{t\in\{p,\frac{n}{2}\}}\bigg{(}\int_{B_{R}}\textit{P}^{t}_{k}(u_{k},u)\,\mathrm{d}x\bigg{)}$
$\leq\displaystyle\sum_{t\in\{p,\frac{n}{2}\}}\bigg{(}\int_{\mathbb{R}^{n}}\textit{P}^{t}_{k}(u_{k},u)\eta\,\mathrm{d}x\bigg{)}$
@@
$\quad-2\sum_{t\in\{p,\frac{n}{2}\}}\Bigg{(}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x\Bigg{)}\quad\text{as}\quad k\to\infty.$ (6.12)
By employing Hölder’s inequality and (6.10), we obtain for $t\in\{p,\frac{n}{2}\}$ that
@@
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\Delta\eta(u_{k}-u)\,\mathrm{d}x=0.$ (6.13)
Similarly, again by using Hölder’s inequality and (6.10), we deduce for $t\in\{p,\frac{n}{2}\}$ that
@@
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\big{(}|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x=0.$ (6.14)
@@
Let $1<s<\frac{n}{\gamma}$ and $r\geq p^{\ast}$ be such that $\frac{1}{\upsilon}+\frac{1}{\upsilon^{\prime}}=1$ for $\upsilon\in\{s,r\}$. Then by using Hölder’s inequality and (6.10), we get
@@
Due to (6.4), passing if necessary to a subsequence, still denoted by $\{u_{k}\}_{k}$, we can assume
@@
Let $\eta,~{}\eta^{\prime}>1$ be such that $\frac{1}{\eta}+\frac{1}{\eta^{\prime}}=1$. From the above inequality, we first fix $m\in(\|u_{k}\|^{\frac{n}{n-2}},\frac{\beta_{\gamma,n}}{\alpha_{0}})$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$ and $\eta>1$ close to $1$ such that $\alpha\eta m<\beta_{\gamma,n}$. It follows from Hölder’s inequality, Theorem 1.3, Corollary 2.6, Corollary 3.1 and (6.10) that
@@
$\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\cfrac{(g(x,u_{k})-g(x,u))(u_{k}-u)\eta}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.15)
Letting $k\to\infty$ in (6) and using (6.13), (6.14) and (6.15), we deduce that $\Delta u_{k}\to\Delta u$ in $L^{\frac{n}{2}}(B_{R})$ as $k\to\infty$ for all $R>0$. Hence, up to a subsequence still denoted by itself such that $\Delta u_{k}\to\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$. It follows that $|\Delta u_{k}|^{t-2}\Delta u_{k}\to|\Delta u|^{t-2}\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$ for $t\in\{p,\frac{n}{2}\}$. Consequently, by using the fact that $\{|\Delta u_{k}|^{t-2}\Delta u_{k}\}_{k}$ is bounded in $L^{\frac{t}{t-1}}(\mathbb{R}^{n})$, we conclude the proof. ∎
@@
Let the hypotheses $(g_{1})-(g_{6})$ be satisfied. Then the functional $J$ satisfies $(PS)_{c}$ condition for any $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 6.4.
@@
$\frac{1}{p}\|\Delta u_{k}\|^{p}_{p}+\frac{2}{n}\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}-\int_{\mathbb{R}^{n}}\frac{G(x,u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\to c\quad\text{as}\quad k\to\infty$ (6.16)
@@
$|\langle{J^{\prime}(u_{k}),\psi}\rangle|\leq\zeta_{k}\|\psi\|,~{}\forall~{}\psi\in E,$ (6.17)
@@
$\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x-\|\Delta u_{k}\|^{p}_{p}-\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}\leq\zeta_{k}\|\ u_{k}\|.$ (6.18)
@@
$\int_{\mathbb{R}^{n}}\frac{|G(x,u_{k})|}{|x|^{\gamma}}\,\mathrm{d}x\leq C\quad\text{and}\quad\int_{\mathbb{R}^{n}}\frac{|g(x,u_{k})u_{k}|}{|x|^{\gamma}}\,\mathrm{d}x\leq C$ (6.19)
for some suitable constant $C>0$. Notice that (6.10) also holds. Due to (1.1), Corollary 3.1, Lemma 2.5 and Hölder’s inequality, one has $\frac{g(x,u)}{|x|^{\gamma}}\in L^{1}_{\textit{loc}}(\mathbb{R}^{n})$. Now by employing a similar strategy developed in de Souza-do Ó [19], we have
$\frac{g(x,u_{k})}{|x|^{\gamma}}\to\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad L^{1}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty.$ (6.20)
Now we claim that
$\frac{G(x,u_{k})}{|x|^{\gamma}}\to\frac{G(x,u)}{|x|^{\gamma}}\quad\text{in}\quad L^{1}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty.$ (6.21)
@@
By using $(g_{4})$ and $(g_{5})$, we can find a constant $K_{0}>0$ such that
$\frac{G(x,u_{k})}{|x|^{\gamma}}\leq K_{0}\frac{|g(x,u_{k})|}{|x|^{\gamma}},~{}\forall~{}x\in\mathbb{R}^{n}.$ (6.22)
Hence we obtain by using (6.10), (6.20), (6.22) and the generalized Lebesgue dominated convergence theorem that
$\lim_{R\to\infty}\lim_{k\to\infty}\int_{B_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$ (6.23)
@@
Due to $(g_{2})$, we can find a constant $C_{1}>0$ such that $|G(x,s)|\leq C_{1}|s|^{\tau}$ for all $|s|\leq K$. It follows from Lemma 2.5 and Theorem 1.7 that

### v3

@@ 2. Preliminary results
Define $\rho=\min\{1,\kappa\}$. Hence, for all $u\in E$ with $\|u\|=\rho$, we have
$J(u)\geq\|u\|^{\frac{n}{2}}\bigg[\frac{2}{n}-C\big(\zeta\|u\|^{\tau-\frac{n}{2}}+\|u\|^{q-\frac{n}{2}}\big)\bigg].$
@@
$\frac{2}{n}-C\big(\zeta\rho^{\tau-\frac{n}{2}}+\rho^{q-\frac{n}{2}}\big)>0.$
@@
$J(u)\geq\rho^{\frac{n}{2}}\bigg[\frac{2}{n}-C\big(\zeta\rho^{\tau-\frac{n}{2}}+\rho^{q-\frac{n}{2}}\big)\bigg]:=\delta>0\quad\text{for}\quad\|u\|=\rho.$
This finishes the proof of the first part of Lemma 8.3. Next, for the proof of the second part, we assume that $u\in E\setminus\{0\}$ and $u\geq 0$ with $\|u\|=1$. Define the map $\Psi\colon[1,+\infty)\to\mathbb{R}$ by $\Psi(s):=s^{-\mu}G(x,su)-G(x,u)$ for all $s\geq 1$ and $u\in\mathbb{R}^{+}:=(0,+\infty)$. By using ($\bf g{4}$), one can easily check that the map $\cdot\mapsto\Psi(\cdot)$ is an increasing function on $1,+\infty)$ and thus, we have $G(x,su)\geq s^{\mu}G(x,u)$ for all $s\geq 1$ and $u\in\mathbb{R}^{+}$. Now, choosing $s\geq 1$ sufficiently large enough, one has
@@
we thank to $\mu>\frac{n}{2}$. The proof is now completed by choosing $e=su$ with $s$ sufficiently large. Hence, the lemma is well-established. ∎
@@
More precisely, we have the following crucial lemma, which describes the range of the mountain pass minimax level $c$.
@@
There exists $\lambda_{0}>0$ sufficiently large enough such that if the hypothesis ($\bf g{6}$) holds for all $\lambda\geq\lambda_{0}$, then the mountain pass minimax level $c$ satisfies
(8.3) $0<c<c_{0}:=\min\Bigg\{\frac{1}{2^{\frac{n}{2}}}\bigg(\frac{2\mu-n}{n\mu}\bigg)^{\frac{n}{2p}}\bigg(\frac{\beta_{\gamma,n}}{\alpha_{0}}\bigg)^{\frac{n-2}{2}},\frac{1}{2^{p}}\bigg(\frac{2\mu-n}{n\mu}\bigg)\bigg(\frac{\beta_{\gamma,n}}{\alpha_{0}}\bigg)^{\frac{(n-2)p}{n}}\Bigg\}.$
@@
Let $\psi\in C^{\infty}_{0}(\mathbb{R}^{n},[0,1])$ be a cut-off function such that $\psi\equiv 1$ in $B_{\frac{1}{2}}$, $\psi\equiv 0$ in $B^{c}_{1}$, $|\nabla\psi(x)|\leq 2$ and $|\Delta\psi(x)|\leq 4$ for all $x\in\mathbb{R}^{n}$. By using ($\bf g{6}$), for any $s\in[0,1]$, we have
@@
Define the map $\pi\colon[0,1]\to E$ by $\pi(s)=s\psi$. Let $\lambda_{1}>0$ be such that for all $\lambda>\lambda_{1}$, we have
@@
It follows that $\pi\in\Gamma$. Moreover, by direct calculations, one has
@@
(8.4) $\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}<\frac{\beta_{\gamma,n}}{\alpha_{0}}.$
@@
By using ($\bf g{4}$) and the above fact, we have
(8.5) $c+o_{k}(1)+o_{k}(1)\|u_{k}\|\geq\bigg(\frac{2}{n}-\frac{1}{\mu}\bigg)\Big(\|\Delta u_{k}\|_{p}^{p}+\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\Big)\quad\text{as}\quad k\to\infty.$
Indeed, if possible, let $\{u_{k}\}_{k}\subset E$ be unbounded in $E$. Thus, we have the following situations.
@@
Moreover, we obtain from (1.9) and (8.5) that
@@
Dividing $\|u_{k}\|^{p}$ on the both the sides and passing $k\to\infty$, we get $0\geq 2^{1-p}\big(\frac{2}{n}-\frac{1}{\mu}\big)>0$, which is a contradiction.
@@
$c+o_{k}(1)+o_{k}(1)\|u_{k}\|_{E}\geq\bigg(\frac{2}{n}-\frac{1}{\mu}\bigg)\|\Delta u_{k}\|_{p}^{p}\quad\text{as}\quad k\to\infty.$
Dividing $\|\Delta u_{k}\|_{p}^{p}$ on the both the sides and passing $k\to\infty$, we get $0\geq\big(\frac{2}{n}-\frac{1}{\mu}\big)>0$, which is again a contradiction.
@@
From the above three cases, we conclude that $\{u_{k}\}_{k}\subset E$ is a bounded sequence in $E$. Thus, by using (8.5), one has the following estimates
(8.6) $\limsup_{k\to\infty}\|\Delta u_{k}\|_{\frac{n}{2}}^{\frac{n}{2}}\leq\bigg(\frac{n\mu}{2\mu-n}\bigg)c\quad\text{and}\quad\limsup_{k\to\infty}\|\Delta u_{k}\|_{p}^{\frac{n}{2}}\leq\bigg(\bigg(\frac{n\mu}{2\mu-n}\bigg)c\bigg)^{\frac{n}{2p}}.$
Recalling the inequality
$(a+b)^{\wp}\leq a^{\wp}+b^{\wp},~\forall~a,b\geq 0\quad\text{and}\quad\wp\in(0,1],$
we obtain from (8.6) that
@@
Notice that $\frac{n\mu}{2\mu-n}>1$ and $\frac{2}{n-2}<\frac{n}{(n-2)p}$. Hence, it follows from the above inequality that
(8.7) $\limsup_{k\to\infty}\|u_{k}\|^{\frac{n}{n-2}}$ $\leq\bigg(2\bigg(\frac{n\mu}{2\mu-n}\bigg)^{\frac{n}{2p}}\bigg)^{\frac{2}{n-2}}\Big(c^{\frac{2}{n-2}}+c^{\frac{n}{(n-2)p}}\Big)$
[…]
Based on (8.3) and (8.7), we immediately conclude the proof. ∎
@@
Let the hypotheses ($\bf g{1}$)($\bf g{4}$) and ($\bf g{6}$) be satisfied. Let $\{u_{k}\}_{k}\subset E$ be an arbitrary $(PS)_{c}$ sequence for $J$ with $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 8.4. Then there exists a subsequence of $\{u_{k}\}_{k}$ still not relabelled and $u\in E$ such that $\Delta u_{k}\to\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$. In addition, there holds
$|\Delta u_{k}|^{t-2}\Delta u_{k}\rightharpoonup|\Delta u|^{t-2}\Delta u\quad\text{in}\quad L^{\frac{t}{t-1}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty\quad\text{for}\quad t\in\Big\{p,\frac{n}{2}\Big\}.$
@@
Let $0<c<c_{0}$ and $\{u_{k}\}_{k}\subset E$ be a $(PS)_{c}$ sequence for $J$. By using Lemma 8.5, $\{u_{k}\}_{k}$ is a bounded sequence in $E$ and (8.4) holds. Hence, we can assume without loss of generality, up to a subsequence, still not relabelled and $u\in E$ such that $u_{k}\rightharpoonup u$ in $E$ as $k\to\infty$. Now, by using Lemma 2.5, Theorem 1.5 and Lemma 2.7, we have
(8.8) $\begin{cases}u_{k}\to u\quad\text{in}\quad L^{\theta}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{for all}\quad\theta\in[1,+\infty);\\ u_{k}\to u\quad\text{in}\quad L^{\varrho}(\mathbb{R}^{n},|x|^{-\gamma}\mathrm{d}x)\quad\text{for all}\quad\varrho\in[p^{\ast},+\infty);\\ \nabla u_{k}\to\nabla u\quad\text{in}\quad L^{t}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{for}\quad t\in\{p,\frac{n}{2}\};\\ u_{k}\to u\quad\text{a.e.~~in}\quad\mathbb{R}^{n}\quad\text{as}\quad k\to\infty.\end{cases}$
@@
(8.9) $\langle{J^{\prime}(u_{k})-J^{\prime}(u),(u_{k}-u)\eta}\rangle=o_{k}(1)\quad\text{as}\quad k\to\infty.$
@@
for any $k\in\mathbb{N}$ and $t\in\{p,\frac{n}{2}\}$. In view of the well-known Simon’s inequality with $n\geq 4$ (for instance, see (8.25)) and (8.9), there exists $c_{\frac{n}{2}}>0$ such that as $k\to\infty$, we have
(8.10) $c^{-1}_{\frac{n}{2}}\int_{B_{R}}$ $|\Delta u_{k}-\Delta u|^{\frac{n}{2}}\,\mathrm{d}x$
$\leq\int_{B_{R}}\textit{P}^{\frac{n}{2}}_{k}(u_{k},u)\,\mathrm{d}x\leq\sum_{t\in\{p,\frac{n}{2}\}}\bigg(\int_{B_{R}}\textit{P}^{t}_{k}(u_{k},u)\,\mathrm{d}x\bigg)$
$\leq\sum_{t\in\{p,\frac{n}{2}\}}\bigg(\int_{\mathbb{R}^{n}}\textit{P}^{t}_{k}(u_{k},u)\eta\,\mathrm{d}x\bigg)$
@@
$\quad-2\sum_{t\in\{p,\frac{n}{2}\}}\Bigg(\int_{\mathbb{R}^{n}}\big(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x\Bigg).$
By employing the Hölder’s inequality and (8.8), we obtain for $t\in\{p,\frac{n}{2}\}$ that
@@
(8.11) $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\big(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\Delta\eta(u_{k}-u)\,\mathrm{d}x=0.$
Similarly, again by using the Hölder’s inequality and (8.8), we deduce for $t\in\{p,\frac{n}{2}\}$ that
@@
(8.12) $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\big(|\Delta u_{k}|^{t-2}\Delta u_{k}-|\Delta u|^{t-2}\Delta u)\nabla\eta\cdot\nabla(u_{k}-u)\,\mathrm{d}x=0.$
@@
Let $1<s<\frac{n}{\gamma}$ and $r\geq p^{\ast}$ be such that $\frac{1}{\upsilon}+\frac{1}{\upsilon^{\prime}}=1$ for $\upsilon\in\{s,r\}$. Then by using the Hölder’s inequality and (8.8), we get
@@
Due to (8.4), passing if necessary to a subsequence, still denoted by itself, we can assume
@@
Let $\eta,~\eta^{\prime}>1$ be such that $\frac{1}{\eta}+\frac{1}{\eta^{\prime}}=1$. From the above inequality, we first fix $m\in\big(\|u_{k}\|^{\frac{n}{n-2}},\frac{\beta_{\gamma,n}}{\alpha_{0}}\big)$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$ and $\eta>1$ close to $1$ such that $\alpha\eta m<\beta_{\gamma,n}$. It follows from the Hölder’s inequality, Theorem 1.1, Corollary 2.6, Corollary 3.1 and (8.8) that
@@
(8.13) $\lim_{k\to\infty}\int_{\mathbb{R}^{n}}\cfrac{(g(x,u_{k})-g(x,u))(u_{k}-u)\eta}{|x|^{\gamma}}\,\mathrm{d}x=0.$
Letting $k\to\infty$ in (8.10) and using (8.11), (8.12) and (8.13), we deduce that $\Delta u_{k}\to\Delta u$ in $L^{\frac{n}{2}}(B_{R})$ as $k\to\infty$ for all $R>0$. Hence, we can assume that up to a subsequence still denoted by itself such that $\Delta u_{k}\to\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$. It follows that $|\Delta u_{k}|^{t-2}\Delta u_{k}\to|\Delta u|^{t-2}\Delta u$ a.e. in $\mathbb{R}^{n}$ as $k\to\infty$ for $t\in\{p,\frac{n}{2}\}$. Consequently, by using the fact that $\{|\Delta u_{k}|^{t-2}\Delta u_{k}\}_{k}$ is bounded in $L^{\frac{t}{t-1}}(\mathbb{R}^{n})$, we conclude the proof. ∎
@@
Let the hypotheses ($\bf g{1}$)($\bf g{6}$) be satisfied. Then the functional $J$ satisfies $(PS)_{c}$ condition for any $0<c<c_{0}$, where $c_{0}$ is defined as in Lemma 8.4.
@@
(8.14) $\frac{1}{p}\|\Delta u_{k}\|^{p}_{p}+\frac{2}{n}\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}-\int_{\mathbb{R}^{n}}\frac{G(x,u_{k})}{|x|^{\gamma}}\,\mathrm{d}x\to c\quad\text{as}\quad k\to\infty$
@@
(8.15) $|\langle{J^{\prime}(u_{k}),\psi}\rangle|\leq\zeta_{k}\|\psi\|,~\forall~\psi\in E,$
@@
(8.16) $\int_{\mathbb{R}^{n}}\frac{g(x,u_{k})u_{k}}{|x|^{\gamma}}\,\mathrm{d}x-\|\Delta u_{k}\|^{p}_{p}-\|\Delta u_{k}\|^{\frac{n}{2}}_{\frac{n}{2}}\leq\zeta_{k}\|\ u_{k}\|.$
@@
(8.17) $\int_{\mathbb{R}^{n}}\frac{|G(x,u_{k})|}{|x|^{\gamma}}\,\mathrm{d}x\leq C\quad\text{and}\quad\int_{\mathbb{R}^{n}}\frac{|g(x,u_{k})u_{k}|}{|x|^{\gamma}}\,\mathrm{d}x\leq C$
for some suitable constant $C>0$. Notice that (8.8) also holds. Due to (1.14), Corollary 3.1, Lemma 2.5 and the Hölder’s inequality, one has $\frac{g(x,u)}{|x|^{\gamma}}\in L^{1}_{\textit{loc}}(\mathbb{R}^{n})$. Now, by employing a similar strategy developed in de Souza-do Ó [19], we have
(8.18) $\frac{g(x,u_{k})}{|x|^{\gamma}}\to\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad L^{1}_{\textit{loc}}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty.$
Next, we claim that
(8.19) $\frac{G(x,u_{k})}{|x|^{\gamma}}\to\frac{G(x,u)}{|x|^{\gamma}}\quad\text{in}\quad L^{1}(\mathbb{R}^{n})\quad\text{as}\quad k\to\infty.$
@@
By using ($\bf g{4}$) and ($\bf g{5}$), we can find a constant $K_{0}>0$ such that
(8.20) $\frac{G(x,u_{k})}{|x|^{\gamma}}\leq K_{0}\frac{|g(x,u_{k})|}{|x|^{\gamma}},~\forall~x\in\mathbb{R}^{n}.$
Hence, we obtain from (8.8), (8.18), (8.20) and the generalized Lebesgue dominated convergence theorem that
(8.21) $\lim_{R\to\infty}\lim_{k\to\infty}\int_{B_{R}}\frac{|G(x,u_{k})-G(x,u)|}{|x|^{\gamma}}\,\mathrm{d}x=0.$
@@
Due to ($\bf g{2}$), we can find a constant $C_{1}>0$ such that $|G(x,s)|\leq C_{1}|s|^{\tau}$ for all $|s|\leq K$. It follows from Theorem 1.5 that